# Optimized 3D Gaussian Splatting using Coarse-to-Fine Image Frequency Modulation

### v1

@@ 4 Results and Evaluation
Implementation Details: All our ablations, graphs and results are reported as the average scores for all held-out test views averaged for all 13 scenes, with the exception of Figure 1 and Figure 5 which report the results averaged for all test views for the bicycle scene. For testing the impact of our module on top of other 3DGS based methods, we first reproduce the software environment and reported results from the authors and then re-run the code with our changes applied using the authors provided training and evaluation scripts for fair comparison and evaluation. We do not change any hyper-parameters or settings for any of the methods except for 3DGS [9] where we change the scaling_lr$=0.01$, densification_interval$=500$ and densify_grad_treshold$=0.0001$. All experiments and comparison were undertaken using a NVidia H200 GPU. We use the step decay function as shown in Figure 4 with a mean blur filter with an initial kernel size of 15 in all our experiments and stop frequency modulation after the first 12,000 iterations. Further ablations of these parameters are presented in Section 4.3. Reconstructed scene models and code to reproduce results will be made publicly available.
@@
Metrics: We use PSNR, SSIM and LPIPS for quantitative evaluation and analysis of the trained models and for reproducing relevant baselines. We follow the established best practice of keeping every 8th frame for a hold-out test set and train all our experiments for 30,000 iterations and comply with the image resolution convention from [9] for using full scale images for DeepBlending [7] and Tanks and Temples [11] and using 2nd and 4th image resolution levels after Colmap processing for MipNeRF360 [1] indoor and outdoor scenes respectively.
4.1 Progressive Growing Impact
Our approach achieves excellent reduction in the number of Gaussians used to represent a scene while maintaining competitive visual quality or improving the visual quality. We achieve a reduction in the number of Gaussians in all datasets for each of the methods as reported in Table 1, which is impressive given that all listed methods aim to reduce the file size and Gaussian count and are already extremely efficient. For some methods, such as MiniSplatting [4] and Compact3DGS [12], the final Gaussian count is not a true representation of the GPU memory requirements as that peaks at various points during training. Opti3DGS also reduces the peak Gaussian load as well as the final Gaussian count for both MiniSplatting[4] and Compact3DGS[12] as reported in Table 1. A reduction in the number of Gaussians during the optimization process also leads to accumulated speedups in terms of optimization time as reported in Table 1 and shown in the Figure 6.
(a) 3DGS* Error Map [PSNR: 25.20, Gaussians: 6,071,666]
(b) Ours Error Map [PSNR: 25.01, Gaussians: 2,107,227]
Figure 5: Pixel level comparison of error maps against the ground truth image. Our approach maintains similar quantitative scores while using fewer Gaussians and results in a decreased pixel wise error. Note the absence of noise like bright points in our results, due to the removal of tiny redundant Gaussians.
We also compare Opti3DGS against the approach used by EAGLES [5] in their coarse-to-fine training strategy where they reduce the image resolution to get a coarse scene representation. However, our approach maintains the full image resolution during the entire image processing pipeline and we show the visual comparison for this in our supplementary material. When isolated, the coarse-to-fine approach from [5], is outperformed by Opti3DGS as reported in Table 2, where we apply both techniques to 3DGS [9] to tease out their impact. Note that we have better reconstruction across the board for all datasets while maintaining equally competitive Gaussian counts. Additionally, if we substitute our frequency modulation approach in the full pipeline from EAGLE [5] as reported in Table 1 there is a positive improvement in results while the Gaussian count is reduced for all reported datasets.
Table 2: We demonstrate the effectiveness of our image filtering approach compared to EAGLES [6]. We use the default image downsizing based approach with starting resize threshold of 0.3 and duration of 0.7 and apply it to 3DGS [9] to isolate its impact. We report our results from ablation on mean filter with a cosine_restart decay schedule with starting kernel size 15.
@@ 4.1 Progressive Growing Impact
Mip-NeRF360 EAGLES [5] (resize 0.3) 0.810 27.44 0.232 2.412
OURS-ablation 0.812 27.45 0.230 2.377
Tanks & Temples EAGLES [5] (resize 0.3) 0.843 23.59 0.191 1.090
OURS-ablation 0.844 23.65 0.188 1.359
Deep Blending EAGLES [5] (resize 0.3) 0.904 29.55 0.247 2.325
OURS-ablation 0.905 29.64 0.245 2.298
4.2 Efficient use of Gaussian Primitives
@@
For 3DGS [9] our approach reduces the Gaussian primitive count by upto 62% but we also makes more efficient use of any given Gaussian budget, meaning it can derive more quality and a better reconstruction given the same number of Gaussian primitives as shown in Table 1. The ADC in 3DGS generally adds redundant and even wasteful Gaussians in some parts of the scene driving up the compute requirements with no gain in quality [18, 12, 19]. Our method reduces the wasteful and redundant use of Gaussians as well and drives the quality higher as shown by the lack of bright noise pattern seen in Figure 5.
[…]
Figure 6: We reduce the GPU memory required during optimization due to a reduced Gaussian primitive load, which also reduces the total optimization time. These results are for 3DGS[9] across the bicycle scenes and are represntive of all scenes tested.
Bicycle Playroom Flowers Counter Train
Ground Truth
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
3DGS[9]
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
3DGS+Ours
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Comp3DGS[12]+Ours
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Mini-Splatting[4]
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Mini-Splatting+Ours
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Reduced-3DGS[19]
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Reduced-3DGS+Ours
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Taming3DGS[15]
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Taming3DGS+Ours
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Figure 7: Qualitative results of Opti3DGS when integrated into different 3DGS-based methods with zoomed in cut-outs.
Figure 8: Ablation Experiments: (a) shows the impact of using different decay rate functions to control the size of the kernel used for image filtering. Combining step decay function with the mean filter performs best and is adopted across all our experiments. Note each point is the average score for that respective dataset. (b) shows the impact of using different initial kernel sizes with the mean filter. Larger starting kernel size reduces the Gaussian count but the quality plateaus.
4.3 Ablations
@@
Modulation Schedule: We experiment with various frequency modulation schedules in Figure 8(a). The results indicate that the step schedule function works better than linear, exponential, or cosine variants across all scenes in the datasets by a significant margin.
@@
Blur Filter: We also test various types of image filtering algorithms including Gaussian, bilinear and mean filtering to determine the best approach. The results, shown in Figure 8(a), indicate that the mean filter performs well on most type of scenes. Surprisingly Gaussian and bilateral filtering perform poorly as they gives more weight to the pixel at the center of the kernel or to edge information in the images. Kernel Size: Holding the step function constant and using a mean filter, we perform experiments to evaluate the impact of the image filtering intensity on the final reconstruction quality. Results, shown in Figure 8(b), indicate that starting from a larger filter improves performance but the trend peaks quickly. We adopt the step decay function with an initial kernel size of 15 in all our experiments for the first 12_000 iterations.
5 Conclusions
@@
Summary: We have presented Opti3DGS, a simple yet highly effective framework that minimizes the number of Gaussian primitives in 3DGS-based pipelines. We demonstrated an average reduction of 62% in Gaussian count, a 40% reduction in the training GPU memory requirements and a 20% reduction in optimization time without sacrificing the visual quality compared to 3DGS [9]. We have conducted an extensive quantitative and qualitative evaluation that has shown Opti3DGS can be seamlessly integrated into a wide range of current 3DGS-based approaches resulting in a positive impact on primitive count in all cases. The contribution of Opti3DGS enables 3DGS-based pipelines and scene reconstructions to be more usable on a range of consumer-grade devices due to reduced primitives and therefore storage, transmission and utilization requirements making the technology more widely accessible.

### v2

@@ 4. Results and Evaluation
Implementation Details: All our ablations, graphs and results are reported as the average scores for all held-out test views averaged for all 13 scenes, with the exception of Figure 1 and Figure 6 which report the results averaged for all test views for the bicycle scene. For testing the impact of our module on top of other 3DGS based methods, we first reproduce the software environment and reported results from the authors and then re-run the code with our changes applied using the authors provided training and evaluation scripts for fair comparison and evaluation. We do not change any hyper-parameters or settings for any of the methods except for 3DGS (Kerbl et al., 2023) where we change the scaling_lr=0.01=0.01, densification_interval=500=500 and densify_grad_treshold=0.0001=0.0001. All experiments and comparison were undertaken using a NVidia H200 GPU. We use the step decay function as shown in Figure 4 with a mean blur filter with an initial kernel size of 15 in all our experiments and stop frequency modulation after the first 12,000 iterations. Further ablations of these parameters are presented in Section 4.3. Reconstructed scene models and code to reproduce results will be made publicly available.
@@
Metrics: We use PSNR, SSIM and LPIPS for quantitative evaluation and analysis of the trained models and for reproducing relevant baselines. We follow the established best practice of keeping every 8th frame for a hold-out test set and train all our experiments for 30,000 iterations and comply with the image resolution convention from (Kerbl et al., 2023) for using full scale images for DeepBlending (Hedman et al., 2021) and Tanks and Temples (Knapitsch et al., 2017) and using 2nd and 4th image resolution levels after Colmap processing for MipNeRF360 (Barron et al., 2022) indoor and outdoor scenes respectively.
4.1. Progressive Growing Impact
Our approach achieves excellent reduction in the number of Gaussians used to represent a scene while maintaining competitive visual quality or improving the visual quality.
Reduced Gaussian Footprint:We achieve a reduction in the number of Gaussians in all datasets for each of the methods as reported in Table 1, which is impressive given that all listed methods aim to reduce the file size and Gaussian count and are already extremely optimized to this end. Averaged across all reported methods for the three datasets, there is a drop of -0.03dB in PSNR for an average drop of 26% in the number of Gaussians. This represents a good trade-off with a - 0.13dB PSNR drop for a 27% drop in #G for Mip-NeRF360, a +0.09dB PSNR gain for a 26% drop in #G for Deep Blending and a - 0.05dB PSNR drop for a 25% drop in #G for Tanks&Temples. The single largest reported PSNR drop in Table 1. of 0.44dB equals a quality loss of 10% while reducing #G by 58% and train time by 16%.
When applied on 3DGS(Kerbl et al., 2023), our method reduces the Gaussian count by up-to 62% as reported in Table 1.
Peak Gaussian Count: For some methods, such as MiniSplatting (Fang and Wang, 2024) and Compact3DGS (Lee et al., 2024), the final Gaussian count is not a true representation of the GPU memory requirements as the Gaussian count peaks at various points during training. Opti3DGS also reduces the peak Gaussian load as well as the final Gaussian count for both MiniSplatting(Fang and Wang, 2024) and Compact3DGS(Lee et al., 2024) as reported in Table 1.
Reduced Optimization Time and GPU Memory: A reduction in the number of Gaussians during the optimization process also leads to accumulated speed-ups in terms of optimization time as reported in Table 1 with an illustrative example visualized in the Figure 5. For 3DGS(Kerbl et al., 2023), our method reduces the optimization time by 20% and the training GPU memory requirements by up-to 40%. Notably, the optimization time for EAGLES(Girish et al., 2024) and Taming3DGS(Mallick et al., 2024)(budget) slightly increases, but this could potentially arise from other hyper-parameters in those methods being optimized for a different training regimes and having a high degree of coupling. Note, we do not change any parameters in any of these methods except for 3DGS. Still, the Gaussian footprint is reliably reduced. The GPU memory required for optimization also decreases proportionally to the reduction in the Gaussian count. An illustrative example of GPU memory reduction is highlighted in Figure 5.
Figure 7. Ablation Experiments: (a) shows the impact of using different decay rate functions to control the size of the kernel used for image filtering. Combining step decay function with the mean filter performs best and is adopted across all our experiments. Note each point is the average score for that respective dataset. (b) shows the impact of using different initial kernel sizes with the mean filter. Larger starting kernel size reduces the Gaussian count but the quality plateaus.
Level-of-Detail Representation: Opti3DGSalso produced different levels-of-details as a natural by-product of our progressive growing approach. We currently only store the last level-of-detail to save disk space. However, it allows a scene to be trained to the maximum level-of-detail before running out of available GPU memory. Figure. 5 and Figure 2 indicate this might be an useful feature and intermediate representations from Opti3DGSmight be more useful than ones obtained naively from 3DGS. We leave further exploration of this idea as future work.
Performance on Scenes with Low Texture: Opti3DGS  gives the highest gains in visual quality on DeepBlending where all scenes include large untextured walls and ceilings with the PSNR improving on all seven reported methods. There is a gain of +0.09dB in PSNR for a 26% drop in #G for Deep Blending and we leave further investigation of this result in context of the SfM induced bias to future work.
Table 2. Compares our progressive filtering approach to the progressive downsampling approach proposed by (Girish et al., 2024). Both are applied on 3DGS(Kerbl et al., 2023) in isolation without any additional changes. We use the default image downsizing based approach with starting resize threshold of 0.3 and duration of 0.7 for (Girish et al., 2024). Our progressive filtering uses the step decay function with mean blur filter and an initial kernel size of 15, which is consistent with the results reported in Table. 1. Results indicate that both approaches have a distinct impact on different datasets.
@@ 4.1. Progressive Growing Impact
Mip-NeRF360 EAGLES(Girish et al., 2024)-Downsizing 0.810 27.44 0.232 2.412
Ours-Filtering 0.808 27.37 0.238 2.195
Tanks & Temples EAGLES(Girish et al., 2024)-Downsizing 0.843 23.59 0.191 1.090
Ours-Filtering 0.843 23.71 0.191 1.229
Deep Blending EAGLES(Girish et al., 2024)-Downsizing 0.904 29.55 0.247 2.325
Ours-Filtering 0.906 29.68 0.246 2.118
4.2. Efficient use of Gaussian Primitives
@@
For 3DGS (Kerbl et al., 2023) our approach reduces the Gaussian primitive count by up-to 62% but it also favours more efficient use of any given Gaussian budget, meaning it can derive more quality and a better reconstruction given the same number of Gaussian primitives. This explains the simultaneous increase of +0.09dB in Deep Blending dataset accompanied by a 26% drop in number of Gaussians used. The ADC in 3DGS generally adds redundant and even wasteful Gaussians in some parts of the scene driving up the compute requirements with no gain in quality (Niedermayr et al., 2024; Lee et al., 2024; Papantonakis et al., 2024). Our method reduces the wasteful and redundant use of Gaussians as well and drives the quality higher as shown by the lack of bright noise pattern seen in Figure 6.
@@
Figure 8. Qualitative results of Opti3DGS when integrated into different 3DGS-based methods with zoomed in cut-outs.
Figure 9. Comparison of our results with 3DGS and respective ground truths. Best viewed zoomed in.
4.3. Ablations
@@
Modulation Schedule: We experiment with various frequency modulation schedules in Figure 4. The results indicate that the step schedule function works better than linear, exponential, or cosine variants across all scenes in the datasets by a significant margin. This is explained by the fact that it allows the ADC to optimize a fixed Gaussian budget rather than optimizing and simultaneously handling new insertions.
@@
Blur Filter: We also test various types of image filtering algorithms including Gaussian, bilinear and mean filtering to determine the best approach. The results, shown in Figure 4, indicate that the mean filter performs well on most type of scenes. Surprisingly Gaussian and bilateral filtering perform poorly, possibly because they give more weight to the pixel at the center of the kernel or to edge information in the images.
Kernel Size: Holding the step function constant and using a mean filter, we perform experiments to evaluate the impact of the image filtering intensity on the final reconstruction quality. Results, shown in Figure 7, indicate that starting from a larger filter improves performance but the trend peaks quickly. We adopt the step decay function with an initial kernel size of 15 in all our experiments for the first 12_000 iterations.
Improvement over EAGLES: EAGLES focuses on storage size and render and train time, starts from 3DGS and uses: (i) MLP based quantization (SH color, opacity and rotation attributes); (ii) influence based Gaussian pruning; (iii) progressive resolution training. Opti3DGS  only uses progressive frequency modulation and 3DGS+Opti3DGS  when compared to the full highly optimized EAGLES pipeline achieves better average Gaussian count by 6.3%, better average Train Time by 4.2% and better average FPS by 43% while there is a minor average PSNR loss of 0.18 dB.
Full-EAGLES with Our Frequency Modulation: ‘EAGLES + Ours’ in Tab. 1 shows that our frequency modulation when plugged into the full EAGLES pipeline as a replacement for its progressive resolution downsampling improves average Gaussian count by 23% with an average PSNR drop of 0.1 dB. Slight loss in quality could be due to EAGLES hyperparameters being optimized for its own resolution modulation and we do not change any of them to ensure a fair and clean comparison.
Core Novelty comparison with EAGLES: EAGLES progressive training module gradually increases image resolution whereas Opti3DGS  only modulates image frequency and when both are applied to 3DGS in isolation our frequency modulation shows a distinct and different impact from resolution downsampling. Our approach reduces Gaussian footprint for two of the three datasets as well as slightly improving the reconstruction quality for two datasets.
5. Conclusions
@@
Summary: We have presented Opti3DGS, a simple yet highly effective framework that minimizes the number of Gaussian primitives in 3DGS-based pipelines. We demonstrated an average reduction of 62% in Gaussian count, a 40% reduction in the training GPU memory requirements and a 20% reduction in optimization time without sacrificing the visual quality compared to 3DGS (Kerbl et al., 2023). We have conducted an extensive quantitative and qualitative evaluation that has shown Opti3DGS can be seamlessly integrated into a wide range of current 3DGS-based approaches resulting in a positive impact on primitive count in all cases. The contribution of Opti3DGS enables 3DGS-based pipelines and scene reconstructions to be more usable on a range of consumer-grade devices due to reduced primitives and therefore storage, transmission and utilization requirements making the technology more widely accessible.